OpenStudy (abbyrose0105):

Simplify 4 over the quantity of negative 2 minus 2i .

OpenStudy (campbell_st):

well multiply it by the conjugate pair to get a rational denominator \[\frac{4}{2 - 2i} \times \frac{2 + 2i}{2 + 2i}\] the denominator will be the difference of 2 squares and you will have to use the identity \[i^2 = -1\] hope it helps

OpenStudy (larseighner):

I think that should be: \[\frac{4}{-2 - 2i} \times \frac{2 + 2i}{2 + 2i}\] but the method is sound and it comes to the same thing since -2 and 2 are both 4 when squared.

OpenStudy (ikram002p):

well no, deffinetly not the same \(4*(2+2i ) \neq 4*(-2+2i )\) conjugate of -2-2i=-2+2i